OpenStudy (anonymous):

simplify ln(e^x * e^3x+4)^(1/2)

OpenStudy (anonymous):

\[\frac{1}{2}(x + \ln(e^{3x} +4))\]

OpenStudy (anonymous):

as i said before ln(e^x)=x so (x + 3x +4)^(1/2) [4(x+1)]^(1/2) 2(x+1)^(1/2)

OpenStudy (anonymous):

Ok Leopold, thanks. I got your answer with a calculator, having a hard time working it out conceptually

OpenStudy (anonymous):

Thats clear to you now? If doesnt I explain more..

OpenStudy (anonymous):

I believe I'm ok for now, many thanks.

OpenStudy (anonymous):

you're welcome